# Learning from Reward-Free Offline Data: A Case for Planning with Latent Dynamics Models

### v3

@@ Abstract
A long-standing goal in AI is to develop agents capable of solving diverse tasks across a range of environments, including those never seen during training. Two dominant paradigms address this challenge: (i) reinforcement learning (RL), which learns policies via trial and error, and (ii) optimal control, which plans actions using a known or learned dynamics model. However, their comparative strengths in the offline setting—where agents must learn from reward-free trajectories—remain underexplored. In this work, we systematically evaluate RL and control-based methods on a suite of navigation tasks, using offline datasets of varying quality. On the RL side, we consider goal-conditioned and zero-shot approaches. On the control side, we train a latent dynamics model using the Joint Embedding Predictive Architecture (JEPA) and employ it for planning. We investigate how factors such as data diversity, trajectory quality, and environment variability influence the performance of these approaches. Our findings reveal that model-free RL excels when abundant, high-quality data is available, while model-based planning demonstrates superior generalization to novel layouts, better trajectory stitching, and greater data efficiency. Notably, planning with a latent dynamics model proves to be a strong approach for handling suboptimal offline data and adapting to diverse environment configurations.
@@ 1 Introduction
Figure 1: Left: We train offline goal-conditioned agents on trajectories collected in a subset of maze layouts (left), and evaluate on held out layouts, observing trajectories shown on the right. Only PLDM solves the task (see Figure˜7 for more). Right: Success rates of tested methods on held-out layouts, as a function of the number of training layouts. Rightmost plot shows success rates of models trained on data from five layouts, evaluated on held-out layouts ranging from those similar to training layouts to out-of-distribution ones. We use map layout edit distance from the training layouts as a measure of distribution shift. PLDM demonstrates the best generalization performance. Results are averaged over 3 seeds, shaded area denotes standard deviation. See Figure˜2 for more details on PLDM.
@@
How can we build a system that performs well on unseen combinations of tasks and environments? One promising approach is to avoid relying on online interactions or expert demonstrations and instead leverage large collections of existing suboptimal trajectories without reward annotations [KPL24, PKL24, Das+20]. In general, two dominant fields offer promising solutions for learning from such data: reinforcement learning and optimal control.
@@
While online reinforcement learning has enabled agents to master complex tasks—from Atari games [Mni13], Go [Sil+16], to controlling real robots [Ope+18]—it demands massive quantities of environment interactions. For instance, [Ope+18] used the equivalent of 100 years of real-time hand manipulation experience to train a robot to reliably handle a Rubik’s cube. To address this inefficiency, offline RL methods [KNL21, Lev+20, EGW05] have been developed to learn behaviors from state–action trajectories with corresponding reward annotations. However, these methods typically train agents for a single task, limiting their reuse in other downstream tasks. To overcome this, recent works have explored learning behaviors from offline reward-free trajectories [Par+24a, TO21, KPL24, PKL24]. This reward-free paradigm is particularly appealing, as it allows agents to learn from suboptimal data and use the learned policy to solve a variety of downstream tasks. For example, a system trained on low-quality robotic interactions with cloth can later generalize to tasks like folding laundry [Bla+24].
@@
Optimal control tackles the challenge differently: instead of learning a policy function via trial and error, it plans actions using a known dynamics model [Ber19, TL05, TES07] to plan out actions. Since real-world dynamics are often hard to specify exactly, many approaches instead learn the model from data [Wat+15, FL17, YCBI19]. This model-based approach has shown generalization in manipulation tasks involving unseen objects [Ebe+18]. Importantly, dynamics models can be trained directly from reward-free offline trajectories, making this a compelling route [Das+20, Ryb+18].
@@
Despite significant advances in RL and optimal control, the role of pre-training data quality on reward-free offline learning remains largely unexplored. Prior work has primarily focused on RL methods trained on data from expert or exploratory policies [Fu+20, Yar+22], without isolating the specific aspects of data quality that influence performance. In this work, we address this gap by systematically evaluating the strengths and limitations of various approaches to learning from reward-free trajectories. We assess how different learning paradigms perform under offline datasets that vary in both quality and quantity. To ground our study, we focus on navigation tasks — an essential aspect of many real-world robotic systems — where spatial reasoning, generalization, and trajectory stitching play a critical role. While this choice excludes other domains such as manipulation, it offers a controlled yet challenging testbed for our comparative analysis.
@@
    We propose two new navigation environments with granular control over the data generation process and generate a total of 23 datasets of varying quality;
@@
    We evaluate methods for learning from offline, reward-free trajectories, drawing from both reinforcement learning and optimal control paradigms. Our analysis systematically assesses their ability to learn from random policy trajectories, stitch together short sequences, train effectively on limited data, and generalize to unseen environments and tasks;
@@
    We present a list of recommendations to help practitioners choose between methods depending on available data and generalization requirements.
To facilitate further research into methods for learning from offline trajectories without rewards, we release code, data, visualizations, and more at our project webpage.
@@ 2 Related Work
Reward-free offline RL refers to learning from offline data that does not contain rewards in a task-agnostic way. The goal is to extract general behaviors from offline data to solve a variety of downstream tasks. One approach uses goal-conditioned RL, with goals sampled in similar manner as in Hindsight Experience Replay [And+17]. [Par+24a] show that this can be applied to learn a goal-conditioned policy using IQL, as well as to learn a hierarchical value function. [Hat+22] propose using a small set of observations corresponding to the solved task to define the task and learn from reward-free data. [Yu+22, Hu+23] propose to use labeled data to train a reward function, than label the reward-free trajectories. Zero-shot methods go beyond goal-reaching from offline data and aim to solve arbitrary tasks specified at test time. For example, HILP [PKL24] proposes learning a distance-preserving representation space such that the distance in that space is proportional to the number of steps between two states, similar to Laplacian representations [WTN18, Wan+21, Wan+22]. Forward-Backward representations [TO21, TRO22] tackle this with an approach akin to successor features [Bar+17].
@@
Optimal Control, similar to RL, tackles the problem of selecting actions in an environment to optimize a given objective (reward for RL, cost for control). Classical optimal control methods typically assume that the transition dynamics of the environment are known [Ber19]. This paradigm has been used to control aircraft, rockets, missiles [Bry96] and humanoid robots [Kui+16, SM09]. When transition dynamics cannot be defined precisely, they can often be learned [Wat+15, TL05a]. Many RL methods approximate dynamic programming in the context of unknown dynamics [Ber12, Sut18]. In this work, we use the term RL to refer to methods that either implicitly or explicitly use rewards information to train a policy function, and optimal control for methods that use a dynamics model and explicitly search for actions that optimize the objective.
@@
The importance of offline data has been highlighted in works such as ExORL [Yar+22], which demonstrates that exploratory RL data enables off-policy algorithms to perform well in offline RL; however, it only compares exploratory vs. task-specific data, without analyzing which data aspects affect performance. [BGB20] investigate the importance of data for offline RL with rewards. Recently proposed OGBench [Par+24] introduces multiple offline datasets for a variety of goal-conditioned tasks; in contrast, we conduct a more fine-grained analysis of how methods perform in top-down navigation under suboptimal data conditions and generalize to new tasks and layouts. [Yan+23] also study generalization of offline GCRL, but focus on reaching out-of-distribution goals. [Ghu+24] study stitching generalization.
@@ 3 The Landscape of Available Methods
In this section, we formally introduce the setting of learning from state-action sequences without reward annotations and overview the available approaches. We also introduce a method we call Planning with a Latent Dynamics Model (PLDM).
@@ 3.1 Problem Setting
We consider a Markov decision process (MDP) ${\mathcal{M}}=({\mathcal{S}},{\mathcal{A}},\mu,p,r)$, where $\mathcal{S}$} is the state space, $\mathcal{A}$} is the action space, $\mu\in{\mathcal{P}}({\mathcal{S}})$ denotes the initial state distribution, $p\in{\mathcal{S}}\times{\mathcal{A}}\rightarrow{\mathcal{S}}$ denotes the transition dynamics (we only consider the deterministic case) and $r\in{\mathcal{S}}\rightarrow{\mathbb{R}}$ denotes the reward function. We work in the offline setting, where we have access to a dataset of state-action sequences $\mathcal{D}$} consisting of transitions $(s_{0},a_{0},s_{1},\ldots,a_{T-1},s_{T})$. We emphasize again that the offline dataset in our setting contains no reward information. The goal is, given $\mathcal{D}$}, to find a policy $\pi\in{\mathcal{S}}\times{\mathcal{Z}}\rightarrow{\mathcal{A}}$, to maximize the cumulative reward $r_{z}$, where $\mathcal{Z}$} is the space of possible task definitions. Our goal is to make the best use of the offline dataset $\mathcal{D}$} to enable the agent to solve a variety of tasks in a given environment with potentially different layouts. During evaluation, unless otherwise specified, the agent is tasked to reach a goal state $s_{g}$, with reward being $r_{g}(s)=\mathbb{I}[s=s_{g}]$, and $\mathcal{Z}$} being equal to $\mathcal{S}$}.
@@ 3.2 Reward-free Offline Reinforcement Learning
Table 1: Road-map of our generalization stress-testing experiments. We test four offline goal-conditioned methods—HIQL, GCIQL, CRL, GCBC; a zero-shot RL method called HILP; and PLDM—a method based on learned latent dynamics and planning. ★★★ denotes good performance in the specified experiment, ★★✩ denotes average performance, and ★✩✩ denotes poor performance. We see that HILP and PLDM are the best-performing methods, with PLDM standing out as the only method that reaches competitive performance in all settings.
@@
Can learn from random policy trajectories (4.5) ★★★ ★✩✩ ★✩✩ ★✩✩ ★✩✩ ★★✩
Can stitch suboptimal trajectories (4.4) ★★★ ★✩✩ ★✩✩ ★✩✩ ★✩✩ ★★✩
In this work, we study methods that solve tasks purely from offline trajectories without reward. Reward-free offline RL methods fall into two categories: goal-conditioned RL and zero-shot methods that treat the task as a latent variable. We evaluate state-of-the-art methods from both categories on goal-reaching, and test zero-shot methods’ abilities to transfer to new tasks. The methods we investigate are:
  • GCIQL [Par+24a]: A goal-conditioned version of Implicit Q-Learning [KNL21], a strong and widely-used method for offline RL;
  • HIQL [Par+24a]: A hierarchical GCRL method which trains two policies: one to generate subgoals, and another one to reach the subgoals. Notably, both policies use the same value function.
  • HILP [PKL24]: A method that learns state representations from the offline data such that the distance in the learned representation space is proportional to the number of steps between two states. A direction-conditioned policy is then learned to be able to move along any specified direction in the latent space.
  • CRL [Eys+22]: A method that uses a contrastive objective to learn compatibility between states and reachable goals. The learned representation is then used to train a goal-conditioned policy.
  • GCBC [Lyn+20, Gho+19]: Goal-Conditioned Behavior Cloning—the simplest baseline for goal-reaching.
@@ 3.3 Planning with a Latent Dynamics Model
The methods listed in Section˜3.2 are model-free, none explicitly model the environment dynamics. Since we do not assume known dynamics as in classical control, we can instead learn a dynamics model from offline data, similar to [NSF20, Per+20], which propose a model-based method for goal-reaching using an image reconstruction objective.
We propose a model-based method named Planning with a Latent Dynamics Model (PLDM), which learns latent dynamics using a reconstruction-free SSL objective and the JEPA architecture [LeC22]. At test time, we plan in the learned latent space to reach goals. We opt for an SSL approach that predicts the latents as opposed to reconstructing the input observations [Haf+23, Fin+16, Zha+19, Ban+18] motivated by findings that reconstruction yields suboptimal features [BL24, Lit+24], while reconstruction-free representation learning works well for control [Shu+20, HWS22]. Appendix˜F provides empirical support: features trained with reconstruction-based methods such as DreamerV3 [Haf+23] underperform in test-time planning.
@@
Encoder: $\quad\hat{z}_{0}=z_{0}=h_{\theta}(s_{0})$ (3.1)
Predictors: $\quad\hat{z}_{t}^{k}=f_{\theta}^{k}(\hat{z}_{t-1},a_{t-1}),\quad k=1,\ldots,K$ (3.2)
where $\hat{z}_{t}^{k}$ is the latent state predicted by predictor $k$ and $z_{t}$ is the encoder output at step $t$. When $K>1$, we train an ensemble of predictors for uncertainty regularization at test time. The training objective involves minimizing the distance between the predicted and encoded latents summed over all timesteps. Given target and predicted latents $Z,\hat{Z}^{k}\in\mathbb{R}^{H\times N\times D}$, where $H\leq T$ is the prediction horizon of the model, $N$ is the batch dimension, and $D$ the feature dimension, the similarity objective between predictions and encodings is:
$\mathcal{L}_{\mathrm{sim}}=\sum_{k=1}^{K}\sum_{t=0}^{H}\frac{1}{N}\sum_{b=0}^{N}\|\hat{Z}^{k}_{t,b}-Z_{t,b}\|^{2}_{2}$ (3.3)
To prevent representation collapse, we use a VICReg-inspired [BPL21] objective, and inverse dynamics modeling [Les+18]. We show a diagram of PLDM in Figure˜2. See Section˜C.1.1 for details.
Figure 3: Left: The Two-Rooms environment. The agent starts at a random location and is tasked with reaching the goal at another randomly sampled location in the other room using 200 steps or less. Observations are $64\times 64$ pixels images. Right: Examples of trajectories in the offline data. Red: each step’s direction is sampled from Von Mises distribution. Blue: each step’s direction is sampled uniformly.
@@
Goal-conditioned planning with PLDM.
$\hat{z}_{0}^{k}=h_{\theta}(s_{0}),\ \hat{z}_{t}^{k}=f_{\theta}^{k}(\hat{z}_{t-1}^{k},a_{t-1}),\quad k=1,\ldots,K$ (3.4)
$C_{\text{goal}}(\mathbf{a},s_{0},s_{g})=\frac{1}{K}\sum_{k=1}^{K}\sum_{t=0}^{H}\|h_{\theta}(s_{g})-f_{\theta}^{k}(\hat{z}_{t}^{k},a_{t})\|$ (3.5)
$C_{\text{unc}}(\mathbf{a},s_{0},s_{g})=\sum_{t=0}^{H}\gamma^{t}\sum_{j=1}^{d}\mathrm{Var}(\{f_{\theta_{k}}(s_{t}^{k},a_{t})_{j}\}_{k=1}^{K})$ (3.6)
$\mathbf{a}^{*}=\operatorname*{argmin}_{\mathbf{a}}\{C_{\text{goal}}(\mathbf{a},s_{0},s_{g})+\beta C_{\text{unc}}(\mathbf{a},s_{0},s_{g})\}$ (3.7)
$C_{\text{goal}}$ is the goal-reaching objective, and $C_{\text{unc}}$ penalizes the model against choosing state-action transitions that deviate from the training distribution, with $\gamma\in[0,1]$ as the temporal discount. This regularization resembles how GCIQL, HIQL, and HILP use expectile regression to learn policies that remain in-distribution under the dataset [KNL22]. See Appendix D for ablations on $C_{\text{unc}}$.
Following the Model Predictive Control framework [ML99], our model re-plans at every $i_{\mathrm{th}}$ interaction with the environment. By default, we use $i=1$ for all experiments. However, the replanning interval $i$ can be extended to improve inference speed at some cost to accuracy (see Appendix˜E). In all our experiments with PLDM, we use MPPI [WAT15] for planning. We note that PLDM does not use rewards, neither explicitly nor implicitly, and should be considered as falling under the optimal control category. We also note that in order to apply PLDM to another task, we do not need to retrain the encoder $h_{\theta}$ and the forward model $f_{\theta}$, we only need to change the definition of the cost in equation Equation˜3.5. We demonstrate this flexibility in Section˜4.6, where we invert the sign of the cost to make the agent avoid the specified state.
4 Stress-Testing Generalization Across Datasets and Tasks
In this section, we conduct thorough experiments to test methods spanning RL and optimal control outlined in Section˜3.2 and Section˜3.3. We evaluate on navigation tasks where the agent is either a point mass (Section˜4.1, Section˜4.8) or quadruped (Section˜4.7). We generate datasets of varying size and quality and test how a specific data type affects a given method. See Table˜1 for overview.

### v4

@@ Abstract
A long-standing goal in AI is to develop agents capable of solving diverse tasks across a range of environments, including those never seen during training. Two dominant paradigms address this challenge: (i) reinforcement learning (RL), which learns policies via trial and error, and (ii) optimal control, which plans actions using a known or learned dynamics model. However, their comparative strengths in the offline setting—where agents must learn from reward-free trajectories—remain underexplored. In this work, we systematically evaluate RL and control-based methods on a suite of navigation tasks, using offline datasets of varying quality. On the RL side, we consider goal-conditioned and zero-shot methods. On the control side, we train a latent dynamics model using the Joint Embedding Predictive Architecture (JEPA) and employ it for planning. We investigate how factors such as data diversity, trajectory quality, and environment variability influence the performance of these approaches. Our results show that model-free RL benefits most from large amounts of high-quality data, whereas model-based planning generalizes better to unseen layouts and is more data-efficient, while achieving trajectory stitching performance comparable to leading model-free methods. Notably, planning with a latent dynamics model proves to be a strong approach for handling suboptimal offline data and adapting to diverse environments.
@@ 1 Introduction
How can we build a system that performs well on unseen combinations of tasks and environments? One promising approach is to avoid relying on online interactions or expert demonstrations, and instead leverage large collections of existing suboptimal trajectories without reward annotations [32, 53, 16]. Broadly, two dominant fields offer promising solutions for learning from such data: reinforcement learning and optimal control.
@@
While online reinforcement learning has enabled agents to master complex tasks—from Atari games [44], Go [60], to controlling real robots [49]—it demands massive quantities of environment interactions. For instance, OpenAI et al. [49] used the equivalent of 100 years of real-time hand manipulation experience to train a robot to reliably handle a Rubik’s cube. To address this inefficiency, offline RL methods [33, 41, 18] have been developed to learn behaviors from state–action trajectories with corresponding reward annotations. However, these methods typically train agents for a single task, limiting their reuse in other downstream tasks. To overcome this, recent work has explored learning behaviors from offline reward-free trajectories [52, 67, 32, 53]. This reward-free paradigm is particularly appealing as it allows agents to learn from suboptimal data and use the learned policy to solve a variety of downstream tasks. For example, a system trained on low-quality robotic interactions with cloth can later generalize to tasks like folding laundry [9].
@@
Optimal control tackles challenge differently: instead of learning a policy function via trial and error, it plans actions using a known dynamics model [7, 64, 63] to plan out actions. Since real-world dynamics are often hard to specify exactly, many approaches instead learn the model from data [71, 20, 77]. This model-based approach has shown generalization in manipulation tasks involving unseen objects [17]. Importantly, dynamics models can be trained directly from reward-free offline trajectories, making this a compelling route [16, 56].
@@
Despite significant advances in RL and optimal control, the role of pre-training data quality on reward-free offline learning remains largely unexplored. Prior work has primarily focused on RL methods trained on data from expert or exploratory policies [22, 76], without isolating the specific aspects of data quality that influence performance. In this work, we address this gap by systematically evaluating the strengths and limitations of various approaches for learning from reward-free trajectories. We assess how different learning paradigms perform under offline datasets that vary in both quality and quantity. To ground our study, we focus on navigation tasks — an essential aspect of many real-world robotic systems — where spatial reasoning, generalization, and trajectory stitching play a critical role. While this choice excludes domains such as manipulation, it offers a controlled yet challenging testbed for our comparative analysis.
@@
    We propose two new navigation environments with granular control over the data generation process, and generate a total of 23 datasets of varying quality;
@@
    We evaluate methods for learning from offline, reward-free trajectories, drawing from both reinforcement learning and optimal control paradigms. Our analysis systematically assesses their ability to learn from random policy trajectories, stitch together short sequences, train effectively on limited data, and generalize to unseen environment layouts and tasks beyond goal-reaching;
@@
    We present a list of guidelines to help practitioners choose between methods depending on available data and generalization requirements.
To facilitate further research into methods for learning from offline trajectories without rewards, we release code, data, environment visualizations, and more at latent-planning.github.io.
@@ 2 Related Work
Reward-free offline RL refers to learning from offline data that does not contain rewards in a task-agnostic way. The goal is to extract general behaviors from offline data to solve a variety of downstream tasks. One approach uses goal-conditioned RL, with goals sampled in similar manner as in Hindsight Experience Replay [1]. Park et al. [52] show that this can be applied to learn a goal-conditioned policy using IQL, as well as to learn a hierarchical value function. Hatch et al. [29] proposes using a small set of observations corresponding to the solved task to define the task and learn from reward-free data. Yu et al. [79], Hu et al. [30] propose to use labeled data to train a reward function, than label the reward-free trajectories. Zero-shot methods go beyond goal-reaching from offline data and aim to solve arbitrary tasks specified at test time. HILP [53] proposes learning a distance-preserving representation space such that the distance in that space is proportional to the number of steps between two states, similar to Laplacian representations [73, 69, 70]. Forward-Backward representations [67, 68] tackle this with an approach akin to successor-features [5].
@@
Optimal Control, similar to RL, tackles the problem of selecting actions in an environment to optimize a given objective (reward for RL, cost for control). Classical optimal control methods typically assumes that the transition dynamics of the environment are known [7]. This paradigm has been used to control aircraft, rockets, missiles [11] and humanoid robots [35, 57]. When the transition dynamics cannot be defined precisely, they can often be learned [71, 65]. Many RL methods approximate dynamic programming in the context of unknown dynamics [6, 62]. In this work, we use the term RL to refer to methods that either implicitly or explicitly use rewards information to train a policy function, and optimal control for methods that use a dynamics model and explicitly search for actions that optimize the objective.
Figure 2: Left: We train offline goal-conditioned agents on trajectories collected in a subset of maze layouts (left), and evaluate on held out layouts, observing trajectories shown on the right. Only PLDM solves the task (see Figure˜8 for more). Right: Success rates of tested methods on held-out layouts, as a function of the number of training layouts. Rightmost plot shows success rates of models trained on data from five layouts, evaluated on held-out layouts ranging from those similar to training layouts to out-of-distribution ones. We use map layout edit distance from the training layouts as a measure of distribution shift. PLDM demonstrates the best generalization performance. Results are averaged over 3 seeds, shaded area denotes standard error. See Figure˜1 for more details on PLDM.
@@
The importance of offline data has been highlighted in works such as ExORL, [76] which demonstrates that exploratory RL data enables off-policy algorithms to perform well in offline RL; however, it only compares exploratory vs. task-specific data, without analyzing which data aspects affect performance. Buckman et al. [12] investigates the data importance for offline RL with rewards. Recently proposed OGBench [51] introduces multiple offline datasets for a variety of goal-conditioned tasks; in contrast, we conduct a more fine-grained analysis of how methods perform in top-down navigation under suboptimal data conditions and generalize to new tasks and layouts. Yang et al. [74] also study generalization of offline GCRL, but focus on reaching out-of-distribution goals. Ghugare et al. [24] study stitching generalization.
@@ 3 The Landscape of Available Methods
In this section, we formally introduce the setting of learning from state-action sequences without reward annotations and overview available approaches. We also introduce a method we call Planning with a Latent Dynamics Model (PLDM).
@@ 3.1 Problem Setting
We consider a Markov decision process (MDP) ${\mathcal{M}}=({\mathcal{S}},{\mathcal{A}},\mu,p,r)$, where $\mathcal{S}$} is the state space, $\mathcal{A}$} is the action space, $\mu\in{\mathcal{P}}({\mathcal{S}})$ denotes the initial state distribution, $p\in{\mathcal{S}}\times{\mathcal{A}}\rightarrow{\mathcal{S}}$ denotes the transition dynamics (we only consider the deterministic case), and $r\in{\mathcal{S}}\rightarrow{\mathbb{R}}$ denotes the reward function. We work in the offline setting, where we have access to a dataset of state-action sequences $\mathcal{D}$} which consists of transitions $(s_{0},a_{0},s_{1},\ldots,a_{T-1},s_{T})$. We emphasize again that the offline dataset in our setting does not contain any reward information. The goal is, given $\mathcal{D}$}, to find a policy $\pi\in{\mathcal{S}}\times{\mathcal{Z}}\rightarrow{\mathcal{A}}$, to maximize cumulative reward $r_{z}$, where $\mathcal{Z}$} is the space of possible task definitions. Our goal is to make the best use of the offline dataset $\mathcal{D}$} to enable the agent to solve a variety of tasks in a given environment with potentially different layouts. During evaluation, unless otherwise specified, the agent is tasked to reach a goal state $s_{g}$, so the reward is defined as $r_{g}(s)=\mathbb{I}[s=s_{g}]$, and $\mathcal{Z}$} is equivalent to $\mathcal{S}$}.
@@ 3.2 Reward-free Offline Reinforcement Learning
Table 1: Road-map of our generalization stress-testing experiments. We test 4 offline goal-conditioned methods - HIQL, GCIQL, CRL, GCBC; a zero-shot RL method HILP, and a learned latent dynamics planning method PLDM. ★★★ denotes good performance in the specified experiment, ★★✩ denotes average performance, and ★✩✩ denotes poor performance. We see that HILP and PLDM are the best-performing methods, with PLDM standing out as the only method that reaches competitive performance in all settings.
@@
Can learn from random policy trajectories (4.5) ★★★ ★✩✩ ★★★ ★✩✩ ★✩✩ ★★✩
Can stitch suboptimal trajectories (4.4) ★★★ ★✩✩ ★★★ ★✩✩ ★✩✩ ★★✩
In this work, we study methods that solve tasks purely from offline trajectories without reward annotations. Reward-free offline RL methods fall into two categories: goal-conditioned RL and zero-shot methods that treat the task as a latent variable. We evaluate state-of-the-art methods from both categories on goal-reaching, and test zero-shot methods on their ability to transfer to new tasks. The methods we investigate are:
  • GCIQL [52] – goal-conditioned version of Implicit Q-Learning [33], a strong and widely-used method for offline RL;
  • HIQL [52] – a hierarchical GCRL method which trains two policies: one to generate subgoals, and another one to reach the subgoals. Notably, both policies use the same value function;
  • HILP [53] – a method that learns state representations from the offline data such that the distance in the learned representation space is proportional to the number of steps between two states. A direction-conditioned policy is then learned to be able to move along any specified direction in the latent space;
  • CRL [19] – uses contrastive learning to learn compatibility between states and possible reachable goals. The learned representation, which has been shown to be directly linked to goal-conditioned Q-function, is then used to train a goal-conditioned policy;
  • GCBC [43, 23] – Goal-Conditioned Behavior Cloning - the simplest baseline for goal-reaching.
@@ 3.3 Planning with a Latent Dynamics Model
The methods in Section˜3.2 are model-free, none explicitly model the environment dynamics. Since we do not assume known dynamics as in classical control, we can instead learn a dynamics model from offline data, similar to [46, 55], which propose a model-based method for goal-reaching using an image reconstruction objective.
We propose a model-based method named Planning with a Latent Dynamics Model (PLDM), which learns latent dynamics using a reconstruction-free SSL objective and the JEPA architecture [39]. At test time, we plan in the learned latent space to reach goals. We opt for an SSL approach that predicts the latents as opposed to reconstructing the input observations [26, 21, 81, 3] motivated by findings that reconstruction yields suboptimal features [2, 42], while reconstruction-free representation learning works well for control [59, 27]. Appendix˜G provides empirical support: features trained with reconstruction-based methods such as DreamerV3 [26] underperform in test-time planning.
@@
$\mathrm{Encoder:}$ $\quad\hat{z_{0}}=z_{0}=h_{\theta}(s_{0})$ (1)
$\mathrm{Predictors:}$ $\quad\hat{z}_{t}^{k}=f_{\theta}^{k}(\hat{z}_{t-1}^{k},a_{t-1}),\forall k\in\{1,\ldots,K\}$ (2)
where $\hat{z}_{t}^{k}$ is the latent state predicted by predictor $k$ and $z_{t}$ is the encoder output at step $t$. When $K>1$, we train an ensemble of predictors for uncertainty regularization at test-time. The training objective involves minimizing the distance between predicted and encoded latents summed over all timesteps. Given target and predicted latents $Z,\hat{Z}^{k}\in\mathbb{R}^{H\times N\times D}$, where $H\leq T$ is the model prediction horizon, $N$ is the batch dimension, and $D$ the feature dimension, the similarity objective between predictions and encodings is:
$\mathcal{L}_{\mathrm{sim}}=\sum_{k=1}^{K}\sum_{t=0}^{H}\frac{1}{N}\sum_{b=0}^{N}\|\hat{Z}^{k}_{t,b}-Z_{t,b}\|^{2}_{2}$ (3)
To prevent representation collapse, we use a VICReg-inspired [4] objective, and inverse dynamics modeling [40]. We show a diagram of PLDM in Figure˜1. See Section˜D.1.1 for details.
Goal-conditioned planning with PLDM.
$\forall k\in\{1,\ldots,K\}:\ \hat{z}_{0}^{k}=z_{0}^{k}=h_{\theta}(s_{0}),\ \hat{z}_{t}^{k}=f_{\theta}^{k}(\hat{z}_{t-1}^{k},a_{t-1})$ (4)
$C_{\text{goal}}(\mathbf{a},s_{0},s_{g})=\frac{1}{K}\sum_{k=1}^{K}\sum_{t=0}^{H}\|h_{\theta}(s_{g})-f_{\theta}^{k}(\hat{z}_{t}^{k},a_{t})\|$ (5)
$C_{\text{uncertainty}}(\mathbf{a},s_{0},s_{g})=\sum_{t=0}^{H}\gamma^{t}\sum_{j=1}^{d}\mathrm{Var}(\{f_{\theta_{k}}(s_{t}^{k},a_{t})_{j}\}_{k=1}^{K})$ (6)
$\mathbf{a}^{*}=\arg\min_{\mathbf{a}}\{C_{\text{goal}}(\mathbf{a},s_{0},s_{g})+\beta C_{\text{uncertainty}}(\mathbf{a},s_{0},s_{g})\}$ (7)
$C_{\text{goal}}$ is the goal-reaching objective and $C_{\text{uncertainty}}$ penalizes the model from choosing state-action transitions that deviate from the training distribution, with $\gamma\in[0,1]$ as the temporal discount. This regularization resembles how GCIQL, HIQL, and HILP use expectile regression to learn policies that remain in-distribution with respect to the dataset [34]. See Appendix E for ablations on $C_{\text{uncertainty}}$.
Following the Model Predictive Control framework [45], PLDM re-plans every $i$ interactions with the environment. By default, we use $i=1$ for all experiments, making PLDM $\sim 4$x slower than the model-free baselines. The replanning interval $i$ can be increased to accelerate MPC with only a minor loss in performance (see Appendix˜F). We use MPPI [72] in all our experiments with planning. We note that PLDM is not using rewards, neither explicitly nor implicitly, and should be considered as an optimal control method. We also note that to apply PLDM to a new task, we do not need to retrain the encoder $h_{\theta}$ and dynamics $f_{\theta}$, we only need to change the cost in Equation˜5. We test this flexibility in Section˜4.6, where we invert the sign of the cost to make the agent avoid a given state.
Figure 3: Left: The Two‐Rooms environment. The agent starts at a random location and is tasked with reaching the goal at another randomly sampled location in the other room using 200 steps or less. Observations are $64\times 64$ pixels images. Right: Examples of trajectories in the offline data. Red: each step’s direction is sampled from Von Mises distribution. Blue: each step’s direction is sampled uniformly.
@@
4 Every Method Can Excel but Few Generalize
In this section, we conduct thorough experiments testing methods spanning RL and optimal control outlined in Section˜3.2 and Section˜3.3. We evaluate on navigation tasks where the agent is either a point mass (Section˜4.1, Section˜4.8) or quadruped (Section˜4.7). We generate datasets of varying size and quality and test how a specific data type affects a given method. See Table˜1 for overview.